Question: Cost of Capital Cloudstreet Ltd is an Australian firm which is publicly-listed on the ASX. The co... Cost of Capital Cloudstreet Ltd is an Australian firm which is publicly-listed on the ASX. The company has a long term target capital structure of 60% Ordinary Equity, 10% Preference Shares, and 30% Debt. All of the shareholders of Cloudstreet are Australian residents for tax purposes. To fund a major expansion Cloudstreet Ltd needs to raise a $120 million in capital from debt and equity markets. CloudstreetLtd’s broker advises that they can sell new corporate bonds to investors for $1030 with a coupon of 6% and a face value of $1,000. Issue costs on this new debt is expected to be 1.5% of face value. The firm can also issue new $100 preference shares which will pay a dividend of $8 and have issue costs of 5%. The company also plans to issue new Ordinary Shares at an issue cost of 3%. The ordinary shares of Cloudstreet are currently trading at $7.50 per share and will pay a dividend of $0.40 this year. Ordinary dividends in Cloudstreet are predicted to grow at a constant rate of 4% pa. a. (i) Calculate how much debt Cloudstreet will need to issue to maintain their target capital structure. (2 marks) (ii) What will be the appropriate cost of debt for Cloudstreet. (8 marks) b. (i) Calculate how much Preference Share equity Cloudstreet will need to issue to maintain their target capital structure. (2 marks) (ii) What will be the appropriate cost of Preference shares for Cloudstreet? (8 marks) c. (i) Calculate how much Ordinary Share equity Cloudstreet will need to issue to maintain their target capital structure. (2 marks) (ii). What will be the appropriate cost of Ordinary Equity shares for Cloudstreet? (8 marks) d. Calculate how the Weighted Average Cost of Capital for Cloudstreet Ltd following the new capital raising. (10 marks) e. Cloudstreet Ltd has a current EBIT of $1.5 million per annum. The CFO approaches the Board and advises them that they have devised a strategy which will lower the company’s cost of capital by a full 1%. How will this change the value of the company? Support your answer using theory and calculations.
Answer #1

Cloudstreeet require $120 million capital for expansion.

Target capital structure =

Equity 60%
Preference share 10%
Debt 30%

a. i) Debt Cloudstreet need to issue to maintain the target capital structure = 30% * 120 = $36 million

Corporate tax rate in Australia =30%

Coupon on the bonds = 6%

a. ii) Cost of debt includes one time issue cost (1.5%) and annual interest cost (6% coupon)

Issue cost of debt = 1.5% *36 = $0.54 million

Annual interest cost on debt= 36 * 6%*(1-30%)= $1.512 million

b.i) Preference share % in target capital structure

Amount of preference shares company need raise to maintain the capital structure = 10% * 120 = $12 million

b.ii) Cost of preference share includes one time issue cost (5%) and annual cost of dividend ($8 per $100 preference share)

Issue cost = 5% *12 = $0.6 million

Annual Dividend cost on preference share= 8/100 *12,000,000= $0.96 million

c. i) Amount of ordinary equity to be raised = 60%*120 = $72 million

c.ii) Cost of ordinary equity will be one time issue cost and cost of dividend

one time issue cost = 3% * 72= $2.16 million

Cost of dividend this year= number of ordinary shares to be issued * dividend per share

= 72,000,000 / 7.5 * 0.4

= 9.6 millions shares * 0.4

Cost of dividend this year =$3.84 million

Dividend on ordinary share will increase 4% every year

d. Weighted average cost of capital

Share price = Dividend * (1+dividend growth) / (cost of equity -dividend growth rate)

Cost of equity =0.4*1.04 /7.5 +4%

= 9.55%

Cost of preference share = 8/100 = 8%

Total cost of capital = 60% * cost of equity + 30% * cost of debt *(1-tax rate) + 10% * cost of preference share

= 60% * 9.55% + 30% * 6% * (1-30%) +10%* 8%

Weighted average cost of capital =7.79%